OpenStudy (anonymous):

2w-16 / 35/w-8 / 25x

OpenStudy (lalaly):

\[\frac{2w-16}{35} \times \frac{25x}{w-8}\]\[\frac{2(w-8)}{35} \times \frac{25x}{w-8}\]\[=\frac{50x}{35} = \frac{10x}{7}\]